Nikki paid Dalton $15.50 for a combination of 15 baseball and football cards. The baseball cards were $1.50 each, the football cards were $0.50 each. How much did Nikki spend on the baseball cards?
Answer by checkley75(3666)

B+F=15 B=15-F
1.50B+.50F=15.50
1.50(15-F)+.50F=15.50
22.50-1.5F+.50F=15.50
-F=15.50-22.50
-F=-7
F=7 IS THE NUMBER OF FOOTBALL CARDS.
B+7=15
B=15-7
B=8 NUMBER OF BASEBALL CARDS.
PROOF
1.50*8+.50*7=15.50
12+3.50=15.50
15.50=15.50